Two traveling sinusoidal waves are described by the wave functions

y1 = 4.80 sin [π(4.10x − 1125t)]
y2 = 4.80 sin [π(4.10x − 1125t − 0.250)]

where x, y1, and y2 are in meters and t is in seconds.
(a) What is the amplitude of the resultant wave function
y1 + y2?
Answer #1

Angle between the phasors = π * 0.250 = π/4 rad

So resultant ampllitude A = sqrt(A1^2+A2^2 + 2A1A2 cos phi )

A = sqrt(4.8^2+4.8^2 + 2*4.8*4.8 cos (π/4) )

= 8.869 m answer